OpenStudy (anonymous):

find d^2y/dx^2 if x^3y^2 = 5 i can find the dy/dx but i have no idea how to find d^2y/dx^2

zepdrix (zepdrix):

Hmmm c:

zepdrix (zepdrix):

\[\large 3x^2y^2+2x^3yy'=0\]\[\large y'=\frac{ -3x^2y^2 }{ 2x^3y }\]\[\large y'=\frac{ -3y }{ 2x }\] Just to make sure you're on the right track, is this what you got for y'? (y' is the same as dy/dx, just in case there's any confusion about that) :)

OpenStudy (anonymous):

yep, i got that. :)

zepdrix (zepdrix):

So from here, taking the derivative with respect to x AGAIN, will give us y'' on the left, anddddd looks like we'll have to apply the quotient rule on the right. :O Need help setting it up?

OpenStudy (anonymous):

yes please!

zepdrix (zepdrix):

\[\large (\frac{ u }{ v })'=\frac{ u'v=uv' }{ v^2 }\]\[\large (\frac{ -3y }{ 2x })'=\frac{ (-3y)'(2x)-(-3y)(2x)' }{ (2x)^2 }\] Ok here is our initial setup, the primes are the terms we need to take a derivative of.

zepdrix (zepdrix):

That part make sense? :D Now we just need to remember, whenever we differentiate a y-term, that a y' will pop out as a result of the chain rule (and because we differentiated WRT x, not y).

OpenStudy (anonymous):

yeah. i get the chain rule, but i still don't get the correct answer... i dont know why...

zepdrix (zepdrix):

Ok you're prolly just missing a simple step, they want the final answer in terms of X and Y ONLY. So see how we end up with a y' somewhere in there? We need to PLUG IN our answer that we got for y' earlier.

OpenStudy (anonymous):

2x * (-3\frac{ dx }{dy}) - (-3) * 2 is that right for the top part?

OpenStudy (anonymous):

oops sorry, i'm new

zepdrix (zepdrix):

hehe ^^

zepdrix (zepdrix):

I dunno, yours is a little hard to read c: lemme know if this is what you're coming up with. \[\large \frac{ -3y'(2x)-(-3y)2 }{ 4x^2 }=\frac{ 6y-6xy' }{ 4x^2 }\]

OpenStudy (anonymous):

\[2x * 3\frac{ dx }{ dy } - 2 * (-3y)\]

OpenStudy (anonymous):

all over 2x^2

zepdrix (zepdrix):

Hmm where'd the minus go on the first term? :O And in your denominator, make sure you're submitting it as (2x)^2, Because 2x^2 is not the same thing :D just in case you're doing this on webassign or something like that.

OpenStudy (anonymous):

but is that right? because the answer is definitely not that. i'm practicing for a test i have tomorrow.

zepdrix (zepdrix):

\[(-3y)'(2x)=(-3\frac{ dy }{ dx })(2x)\] The minus disappeared from your first term here, careful nowww :O

OpenStudy (anonymous):

oops, yeah. i missed that.

zepdrix (zepdrix):

Earlier we had found the FIRST derivative to be this: \[\large \frac{ dy }{ dx }=\frac{ -3y }{ 2x }\] So to clean up our SECOND derivative, we should plug this in for dy/dx.

OpenStudy (anonymous):

so is what i have the answer then?

zepdrix (zepdrix):

I think this is the answer you want to try and get.

zepdrix (zepdrix):

requires quite a bit of simplification :O

OpenStudy (anonymous):

\[- \frac{ 3 }{ 2 } \left( x \frac{ dy }{ dx } -y \right)\] then all the stuff inside the brakets over x^2 that's the answer given

zepdrix (zepdrix):

Oh they didn't change the dy/dx to x's and y's? Oh i see, they were taking it easy on you XD interesting.. i went too far with that, my bad.

zepdrix (zepdrix):

sec ill take another look :3

OpenStudy (anonymous):

okie. thanks by the way for the help

zepdrix (zepdrix):

\[\huge \frac{ 6y-6x \frac{ dy }{ dx } }{ 4x^2 }=\frac{ -6(x \frac{ dy }{ dx }-y) }{ 4(x^2) }\]

zepdrix (zepdrix):

\[\huge =\frac{ -3 }{ 2 }(\frac{ x \frac{ dy }{ dx } -y}{ x^2 })\]

OpenStudy (anonymous):

?!?!

zepdrix (zepdrix):

Understand the steps I took? :O it's a lil tricky, bunch of stuff going on. A lil bit of factoring and such.

OpenStudy (anonymous):

i'm trying to follow it. i get some of it... i don't know why all the stuff gets moved around though...

zepdrix (zepdrix):

Were you able to get it to this form yet? \[\huge \frac{ 6y-6x \frac{ dy }{ dx } }{ (2x)^2 }\]

OpenStudy (anonymous):

\[-6x \frac{ dy }{ dx } + 6y\] that's what i got so far... so yeah i guess it's the same if i move that 6y over to the other side...

zepdrix (zepdrix):

Yah subtraction looks prettier if you write it on the right ^^ hehe

OpenStudy (anonymous):

and i get how you get 4x^2 on the bottom. but what i'm now confused with is why the answer is broken up the way it is...

zepdrix (zepdrix):

There is no reason for it, the way they broke it up is quite unnecessary. If you take a test, your teacher is very unlikely to take off any points for not simplifying it to that exact point. You should probably notice that the top and bottom can have a 2 divided out of them. But that's the only simplification necessary. The rest was just your book being silly (or did teacher give you this problem in class?).

OpenStudy (anonymous):

this was a practice test from the teacher. and that was the solution. i also have another questoin. how did the -6y get to just - y?

zepdrix (zepdrix):

 Understand how factoring works? :)